OpenStudy (anonymous):

can anyone please solve this What is the unit's digit in the product of (7^71*6^59*3^65)????

OpenStudy (anonymous):

you need to solve : (7^71*6^59*3^65) mod 10

OpenStudy (anonymous):

\( 7^{71} = 7(7^2)^{35} = 7(-1)^{35} = -7 = 3 \mod 10 \)

OpenStudy (anonymous):

similarly find the other remainders

OpenStudy (anonymous):

3^x can have only three digits in the unit place (leaving the case when x=0) 3,9,1 so for 3^65 you do 65 mod 3 which give 2 so unit place will be 9 for 3^65 similarly 6^x has only 1 value in the unit place so 6^59 will have 6 in the unit place for 7^x we have 4 values 7,9,3,1 in the unit place so we will find 71 mod 4 which is 3 so 7^71 will have 3 in the unit place Now we will get 3*6*9 =162 hence 2 will be in the unit place

OpenStudy (anonymous):

but the answer to this question is 4

OpenStudy (zpupster):

Unit's digit of (3^65)*(6^59)*(7^71) = (Unit's digit of 3^65)*(Unit's digit of 6^59)*(Unit's digit of 7^71) Any power of 6 ends with 6. Hence, unit's digit 6^59 = 6 Now for powers of 3, Unit's digit of 3^(Multiple of 4) = 1 Unit's digit of 3^(Multiple of 4 + 1) = 3 Unit's digit of 3^(Multiple of 4 + 2) = 9 Unit's digit of 3^(Multiple of 4 + 3) = 7 Hence, unit's digit of 3^65 = unit's digit of 3^(4*16 + 1) = 3 Now for powers of 7, Unit's digit of 7^(Multiple of 4) = 1 Unit's digit of 7^(Multiple of 4 + 1) = 7 Unit's digit of 7^(Multiple of 4 + 2) = 9 Unit's digit of 7^(Multiple of 4 + 3) = 3 Hence, unit's digit of 7^71 = unit's digit of 7^(4*17 + 3) = 3 Hence, unit's digit of (3^65)*(6^59)*(7^71) = unit's digit of 3*6*3 = unit's digit of 54 = 4

OpenStudy (anonymous):

ok got it thanks a lot

OpenStudy (zpupster):

you are welcome!!

OpenStudy (anonymous):

K I made a mistake with 3. Sorry

OpenStudy (anonymous):

@beccaboo022a can you help me understand your method please

OpenStudy (anonymous):

thanks for helping:)

OpenStudy (anonymous):

\(7^{71} = 7(7^2)^{35} \)

OpenStudy (anonymous):

@thisSucks

OpenStudy (anonymous):

k then?

OpenStudy (anonymous):

\(7^2 = 49\) leaves a remainder of -1 when divided by 10

OpenStudy (anonymous):

so, \(7^{71} = 7(7^2)^{35} = 7(49)^{35} = 7(-1)^{35} \mod 10 \)

OpenStudy (anonymous):

see if that makes sense

OpenStudy (anonymous):

I followed till the point 72=49 leaves a remainder of -1 when divided by 10 But not sure how can we write 7(49)35=7(−1)35 mod 10

OpenStudy (anonymous):

we use this property if mods :- \(a = b \mod n \) \(\implies \) \(a^k = b^k \mod n \)

OpenStudy (anonymous):

*of

OpenStudy (anonymous):

\(49 = -1 \mod 10\) \(\implies\) \(49^k = (-1)^k \mod 10\)

OpenStudy (anonymous):

I have not learned this property. Also I learned that mod always gives +ve number for positive numbers.

OpenStudy (anonymous):

okay, thats just false. -1 is same as 9 in mod 10

OpenStudy (anonymous):

-11 = -1 = 9 in mod 10

OpenStudy (anonymous):

mod is defined for all integers

OpenStudy (anonymous):

Thanks for the effort but still I am not clear. even when I try it on google calculator -1 mod 10 gives 9

OpenStudy (anonymous):

yes, so 9 mod 10 = -1 mod 10

OpenStudy (anonymous):

Thanks! Can you share any online proof for a=b mod n I don't want to waste your time

OpenStudy (anonymous):

i have time, i can try the proof here. its simple, wont take long time

OpenStudy (anonymous):

Sure!

OpenStudy (anonymous):

can you state the proof here. I will have a look and if any doubt I will ask you

OpenStudy (anonymous):

first a simple definition for mod :- \(a \equiv b \mod n\) means, \(n | a-b\)

OpenStudy (anonymous):

next few trivial results :- 1) \(a \equiv a \mod n\) 2) if \(a = b \mod n \), then \(b \equiv a \mod n \) 3) if \(a = b \mod n \) and \(c = d \mod n \), then \(a+c = b+d \mod n \) and \(ac = bd \mod n \)

OpenStudy (anonymous):

next we can conlcude the proof, if u are okay wid above results :)

OpenStudy (anonymous):

K let me comprehend

OpenStudy (anonymous):

yahh

OpenStudy (anonymous):

to make sense of above results : for each result, just verify if the definition of mod holds or not.

OpenStudy (anonymous):

I feel so stupid...ugh But not sure how 3) point works. It holds true for values though

OpenStudy (anonymous):

okay lets prove 3rd result, before moving to our actual proof :)

OpenStudy (anonymous):

No No not required I realized it how it works :)

OpenStudy (anonymous):

good :) next we prove below using induction :- to prove : if \(a \equiv b \mod n\), then \(a^k \equiv b^k \mod n \)

OpenStudy (anonymous):

clearly, the statement holds when \(k = 1\)

OpenStudy (anonymous):

Yeah I think I can prove it with induction. But I never get a good understanding when proofs are done via induction

OpenStudy (anonymous):

oh actually induction proofs are easy to *see* than analytical proofs

OpenStudy (anonymous):

as for: Unit's digit of 3^(Multiple of 4) = 1 < same Unit's digit of 3^(Multiple of 4 + 1) = 3 Unit's digit of 3^(Multiple of 4 + 2) = 9 Unit's digit of 3^(Multiple of 4 + 3) = 7 so, the unit's digit of 3^65 = unit's digit of 3^(4*16 + 1)=3 and when the powers of 7... the unit's: Unit's digit of 7^(Multiple of 4) = 1 < same Unit's digit of 7^(Multiple of 4 + 1) = 7 < Unit's digit of 7^(Multiple of 4 + 2) = 9 Unit's digit of 7^(Multiple of 4 + 3) = 3 < .....

OpenStudy (anonymous):

ever looked at proof for induction method ?

OpenStudy (anonymous):

I have proofed it using property 3.

OpenStudy (anonymous):

if u did, then u wud love induction proofs

OpenStudy (anonymous):

Yes.... property3 is needed for induction here...

OpenStudy (anonymous):

I mean there is a proof, which proves that induction method works.

OpenStudy (anonymous):

But somehow I don't *see* things with induction but that's OK. But still I say lets divide the step 2 in two parts and say -1 mod 10 is equivalent to 49 mod 10

OpenStudy (anonymous):

I would love to see the proof that says induction method works :)

OpenStudy (anonymous):

This is a cool method. I liked it Thanks :) K and help me with one more thing...How can I reward you for all your effort :) I am in debt now lol

OpenStudy (anonymous):

lol im in debt too:) proof of induction method is simple :- see if u can get along wid this : Theorem 1.2 First Principle of Finite Induction. Let S be a set of positive integers with the following properties: (a) The integer 1 belongs to S. (b) Whenever the integer k is in S, the next integer k + 1 must also be in S. \(\text{Then S is the set of all positive integers.} \)

OpenStudy (anonymous):

thats the statement for 'induction method'. we can prove it elegantly wid a small trivial argument.

OpenStudy (anonymous):

I know the statement. and It makes sense but how can you say that proving that k+1 is the *sufficient* condition

OpenStudy (anonymous):

exactly, the argument involves proving that, 'proving k+1' is sufficient :)

OpenStudy (anonymous):

BTW you are really smart. Do you actually remember the definitions also?

OpenStudy (anonymous):

Ohh yes I get it now, A bit thoughtful Thanks :)

OpenStudy (anonymous):

(a) The integer 1 belongs to S. so, 1 is in set. (b) Whenever the integer k is in S, the next integer k + 1 must also be in S. assuming k is in set, IF WE COULD PROVE that k+1 also is in set --- thats same as proving that the set contains all integers >=1

OpenStudy (anonymous):

Yeah I get it now! Thank you so much :)

OpenStudy (anonymous):

i have another ques can u ppl pis solve it now

OpenStudy (anonymous):

I see np :) thanks for the kind words lol you're super smart xD

OpenStudy (anonymous):

yah losan you may close this and ask it in a thread :)

OpenStudy (anonymous):

*new thread

OpenStudy (anonymous):

Yeah let me try this one with all the extra knowledge I gained today

OpenStudy (anonymous):

ok